OpenStudy (anonymous):

solve : 3^4x+7 = 4^2x+3.....solve for x

OpenStudy (amistre64):

this one?

OpenStudy (amistre64):

are those all in the exponent? (..) would help it they are

OpenStudy (amistre64):

3^(4x+7) = 4^(2x+3) log3 (3^(4x+7) = 4^(2x+3)) 4x+7 = (2x+3) log3(4) 4x+7 ----- = log3(4) 2x+3 hmmm....

OpenStudy (anonymous):

what does the log do?

OpenStudy (amistre64):

\[b^{n+m}=b^n*b^m\]might help 3^(4x+7) = 4^(2x+3) 3^(4x) * 3^7 = 4^(2x)* 4^3 3^(4x) * 3^7 = 4^(2x)* 4^3 3^(4x) = 4^(2x) * N 3^(4x) ------ = N 4^(2x) log helps to undo exponent bases

OpenStudy (anonymous):

ok thanks a lot!!! you sound like you know a lot about this stuff...im in precalc and these are things we havent relee learned yet

OpenStudy (amistre64):

log4 (3^(4x) = 4^(2x) * N) 4x log4(3) = 2x + log4(N) 4x log4(3) - 2x= log4(N) x(4 log4(3) - 2)= log4(N) log4(N) x= -------------- log4(3^4) - 2 log4(4^3/3^7) x= -------------- log4(3^4) - 2 3-7 log4(3) x= ------------- 4 log4(3) - 2 maybe, might wanna chk with the wolf on that tho